ECONOMICS 504 MATHEMATICAL STATISTICS ROBIN SICKLES FALL 1999

TEXTS: Spanos, A.: Statistical Foundations of Econometric Modelling, 1986, (S) Greene, W. H.: Econometric Analysis, 1998, Fourth Edition, (G) Amemiya, T.: Introduction to Statistics and Econometrics, 1994, (A) Dudewicz, E. J., and S. N. Mishra, Modern Mathematical Statistics, 1988,

(DM)

EXAMINATIONS

Three take home problem sets. Inclass openbook midterm and final exam. Final exam will be 30% of the final grade. Homework exercises and midterm each will make up 35% of final grade.

COURSE OUTLINE:

A. Probability Theory

1. The Calculus of Probability S- 3.1-3.3; G- 3.1-3.2.2, 3.6-3.7.3; A- 2.5; DM-Chp. 1-2.

2. Mathematical Expectations and Moments of Random Variables S- 4.1-4.3, 4.5, 5.1, 5.3, 5.4.; G- 3.3, 3.5-3.7; A- 4;DM-Chp 5.

3. Special Distributions S- 4.4, 5.2, Appendix 6.1; G- 3.4, 3.8, 3.10; A- 5;DM-Chp. 3-4.5

4. Probability Transformations:

S- 6.1-6.3, 7; DM-Chp.4.5-4.8

5. Asymptotic Theory and Limiting Distributions:

S- 9.1-9.3, 10.1-10.2; G- 4.4.1-4.4.2; A- 6.1-6.2;DM-Chp. 6.

B. Statistical Inference and Regression

1. Sampling Distributions:

S- 11.1-11.2, 11.5; G- 4.1-4.2, 4.4.1.

2. Estimation S- 12.1-12.3, 13.1-13.3; G- 4.3, 4.4.6, 4.5-4.6; A- 7, 8; DM-Chp. 7-8.

3. Testing for Statistical Hypothesis:

S- 11.3-11.4, 14.1-14.6; G- 4.7; A- 9.1-9.2; DM-Chp. 9.

4. Regression and Correlation:

G- 5.1-5.6, 6.1-6.3; A- 10.2-10.3; DM-Chp. 14

5. Bayesian Methods:

A- 8.3.